how many angular nodes are present in 5dz^2.how can we calculate it?

Grade:11

1 Answers

abinayaa
5 years ago

For a given orbital, there are two types of nodes :

The number of angular nodes = l

The number of radial nodes = (n - l- 1)

Total number of nodes = n - 1

Where:

n = Principal quantum number

l = Azimuthal quantum number

So, for the 5 d orbital, n = 5 and l = 2.

Number of angular nodes = 2

Number of radial nodes = 5-2-1=2

And, total number of nodes = 3